Shift summary for 10/12/99 Owl Shift

Julie Whitmore (jaws@fnal.gov)
Tue, 12 Oct 1999 07:04:09 -0500 (CDT)

Shift Summary for Owl shift, 12/Oct/99

On Shift:Jianbo, Julie

Hours of beam available at all intensities: 6.25
Hours of beam available at nominal intensity: 6.25
Hours of E799 data (special runs not included): 6.08
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Tape change and dropped interlock.

========================================================================
Runs ended this shift:
Run 14822 - 12.8M events - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

UP*61-66,UPS21-22

========================================================================
Accesses during shift: Search and secure.

========================================================================
Overview of shift:
Relatively smooth data taking until interlock spontaneously drops.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14822 00:00 Take over BEAM_799 run from evening crew.
SDAQ Epicure button red due to low readback
on NM2SR. This is a known problem.

198 00:45 TRD8 trips off. Page Sasha.

214 01:05 TRD7 trips off while SashaL is working
on TRD8. Resets TRD8.

235 01:35 SashaL resets TRD7 but it continues
to have problems. In order to
keep TRD7 from tripping off during the
owl shift, he reduces the voltages by 60V.
He will try to restore voltages to nominal
tomorrow morning, when he can monitor them.

02:55 Tapes full. Run ends.

03:00 Load ADGN and DCT0 constants.

14823 03:05 Start run with fresh tapes UP*67-69,UPS23.

03:10 Monitor stats display dies. Restart.

25 03:35 Scaler DAQ dead for 2.9sec. Beam intensity
up to 9.5E12. Drops back after a spill.

30-34 03:45 No beam.

42 04:02 KTEVANA errors. KFREND.

45 04:05 HCC inefficient. Reload Et.

52 04:16 Beam gone. MCR initially thinks it is a
rad trip but it turns out that our interlocks
have dropped again. Need a Search&Secure.

04:45 S&S finished. Suspect interlock dropped by
beetles on south door interlock sensor. Bugs
removed by ops and noted in MCR online log.

81 04:51 Beam returns.

05:19 Onsplit request: UPE004 on ktev2.

05:50 Onsplit request: UPK008 on ktev4.

06:25 Onsplit request: UPX005 on ktev3.

150-155 06:25 No beam.

172-177 06:56 No beam. Hand off to day crew.

========================================================================
Follow up Notes on Previous Shift

Installed DCT0 constants.

========================================================================
Items to Follow up on Next Shift

Keep an eye on TRDs. Page SashaL if they trip off.

Known problems
--------------
SDAQ: Epicure button red - NM2SR readback low.

Online histos: Banana 1X-3 (sometimes 1X-2), RC10 timing,
DC planes 9&10 inefficiency in beam region,
DC4Y intrinsic efficiency low, DC3X high SOD rms,
CsI 961 comparator errors, CsI 2310 7% gain drift.